A bus and a train leave a station at the same time.the bus travels soouth at 55 mph. the van heads west at 59 mph. There two way communication has a 250 mile range, how long could they drive before they were out of range with each other in hours and minutes?
Answer by checkley77(12844)

THIS IS A RIGHT TRIANGLE WITH SIDES=55T, 59T & 250.
(55T)^2+(59T)^2=250^2
3,025T^2+3,481T^2=62,500
6,506T^2=62,500
T^2=62,500/6,506
T^2=9.6065
T=SQRT9.6065
T=3.1 HOURS OR 3 HRS. 6 MIN.
PROOF:
(55*3.1)^2+(59*3.1)^2=62,500
170,5^2+182.9^2=62,500
29,070.25+33,452.41=62,500
62,500~62,500